# Interpretable and Steerable Concept Bottleneck Sparse Autoencoders

### v1

@@ Appendix A Implementation Details
Here, $c_{k}$ is the $k^{\text{th}}$ concept prediction for the current mini-batch and $\hat{y}_{k}$ is the zero-shot CLIP prediction for concept $k$ with the same mini-batch. Following [oikarinen2023labelfree], we also normalize both vectors $c_{k},\hat{y}_{k}\;\forall\;k$ before raising them to the third power (element-wise) and computing the cosine similarity. The third power is used to make the loss more sensitive to highly activating inputs. And we minimize the negative similarity which is equivalent to maximizing the similarity.
@@
Appendix B Experiments
B.1 Experimental Setup
Downstream model details. We experiment with SAEs/CB-SAEs trained on vision encoders for downstream models like LLaVA [liu2023visual] and UnCLIP [Rombach_2022_CVPR]. LLaVA models are large vision-language models that take an image and a text prompt as input and output a text-based answer (Fig. 2A, main paper). Specifically, we used LLaVA-1.5-7B [liu2024improved] which uses a CLIP-ViT-L-14-336 [radford2021learning] vision encoder, a 2-layer MLP projector (not shown in Fig. 2A for simplicity), and an instruction-finetuned Vicuna-7B LLM [vicuna2023]. We also use LLaVA-MORE [cocchi2025llava] with DINOv2-Large [oquab2024dinov2] vision encoder, a 2-layer MLP projector, and an instruction-finetuned Gemma2-9B LLM [team2024gemma]. On the other hand, UnCLIP is an image-to-image generative model that uses a CLIP-ViT-L [radford2021learning] vision encoder and a finetuned Stable Diffusion 2.1 [Rombach_2022_CVPR] as the image generator (Fig. 2B, main paper).
Miscellaneous details. We implement our CB-SAE in PyTorch [paszke2019pytorch] building on the SAE codebase from pach2025sparse. Following the baseline SAE training [pach2025sparse], we train the CB-SAE for 110k iterations with batch size 4096 and learning rate 2e-4 on a single 80GB Nvidia H100 GPU.
B.2 Interpretability vs Steerability in SAEs
@@
None (SAE baseline) [pach2025sparse] 74.07 0.154 0.517 0.198 0.203
@@
CLIP [radford2021learning] ViT-B-16 0.198 0.307
CLIP [radford2021learning] ViT-L-14-336 0.154 0.244
SigLIP [zhai2023sigmoid] ViT-SO400M-14-384 0.189 0.289
SigLIP2 [tschannen2025siglip] ViT-gopt-16-384 0.188 0.290
SigLIP2 [tschannen2025siglip] ViT-SO400M-16-384 0.176 0.272
DFN [fang2024data] ViT-H-14-378 0.220 0.347
PE-core [bolya2025perception] BigG-14-448 0.207 0.312
B.3 Extended Analysis of our CB-SAE
@@
Sensitivity to scores used for SAE pruning. We extend our sensitivity analysis from Fig. 5A (main paper) in Table 3 to additionally include monosemanticity score [pach2025sparse] (interpretability evaluation) and zero-shot ImageNet-1k accuracy (reconstruction evaluation) when using the SAE/CB-SAE reconstructed latents. We observe that using either the interpretability score or both scores yields similar reconstruction as the baseline SAE, while steerability-based pruning leads to significantly worse reconstruction. Similarly, using either the interpretability score or both scores improves the monosemanticity significantly w.r.t. the baseline, while steerability-based pruning provides only a marginal gain over the baseline.
@@
Sensitivity to CLIP model in interpretability evaluation. We evaluate the sensitivity of our interpretability evaluation with CLIP-Dissect by varying the CLIP-like model used, in Table 4. While our evaluation used a stronger CLIP-ViT-L-14-336 [radford2021learning] model w.r.t. the smaller CLIP-ViT-B-16 used for training the CB-SAE, we now evaluate with even stronger models including SigLIP [zhai2023sigmoid], SigLIP2 [tschannen2025siglip], Data Filtering Networks (DFN) [fang2024data] and Perception Encoder (PE) [bolya2025perception]. Across all CLIP-like models, our CB-SAE achieves consistent gains over the baseline SAE for LLaVA with CLIP-ViT-L encoder, validating that our choice of CLIP-like model for interpretability score does not affect our evaluation.
@@
B.4 Extended Qualitative Results

### v2

@@ Appendix B Implementation Details
Here, $c_{k}$ is the $k^{\text{th}}$ concept prediction for the current mini-batch and $\hat{y}_{k}$ is the zero-shot CLIP prediction for concept $k$ with the same mini-batch. Following [34], we also normalize both vectors $c_{k},\hat{y}_{k}\;\forall\;k$ before raising them to the third power (element-wise) and computing the cosine similarity. The third power is used to make the loss more sensitive to highly activating inputs. And we minimize the negative similarity which is equivalent to maximizing the similarity.
@@
Appendix C Experiments
C.1 Experimental Setup
Downstream model details. We experiment with SAEs/CB-SAEs trained on vision encoders for downstream models like LLaVA [29] and UnCLIP [46]. LLaVA models are large vision-language models that take an image and a text prompt as input and output a text-based answer (Fig. 2A, main paper). Specifically, we used LLaVA-1.5-7B [28] which uses a CLIP-ViT-L-14-336 [42] vision encoder, a 2-layer MLP projector (not shown in Fig. 2A for simplicity), and an instruction-finetuned Vicuna-7B LLM [11]. We also use LLaVA-MORE [12] with DINOv2-Large [36] vision encoder, a 2-layer MLP projector, and an instruction-finetuned Gemma2-9B LLM [52]. On the other hand, UnCLIP is an image-to-image generative model that uses a CLIP-ViT-L [42] vision encoder and a finetuned Stable Diffusion 2.1 [46] as the image generator (Fig. 2B, main paper).
Miscellaneous details. We implement our CB-SAE in PyTorch [39] building on the SAE codebase from Pach et al. [37]. Following the baseline SAE training [37], we train the CB-SAE for 110k iterations with batch size 4096 and learning rate 2e-4 on a single 80GB Nvidia H100 GPU.
Table 3: Sensitivity to type of SAE.
SAE type Interpretability Steerability
CD MS Dead Neurons Unit-Vec White Image
Top-$k$ SAE 0.162 0.548 52965 / 65536 0.228 0.241
Batch Top-$k$ SAE 0.158 0.540 56899 / 65536 0.226 0.231
Matryoshka SAE 0.154 0.517 4 / 65536 0.198 0.203
Top-$k$ CB-SAE 0.264 0.556 0 / 32167 0.315 0.317
Batch Top-$k$ CB-SAE 0.265 0.564 0 / 32162 0.307 0.299
Matryoshka CB-SAE 0.244 0.556 4 / 32169 0.261 0.250
Table 4: Ablation to quantify the usefulness of SAE in CB-SAE.
Interpretability Steerability
CLIP-Dissect score Unit-Vec White Image
SAE [37] 0.154 0.198 0.203
CB-AE (w/o SAE) 0.308 0.238 0.232
CB-SAE (Ours) 0.244 0.261 0.250
C.2 Interpretability vs Steerability in SAEs
@@
None (SAE baseline) [37] 74.07 0.154 0.517 0.198 0.203
@@
CLIP [42] ViT-B-16 0.198 0.307
CLIP [42] ViT-L-14-336 0.154 0.244
SigLIP [64] ViT-SO400M-14-384 0.189 0.289
SigLIP2 [54] ViT-gopt-16-384 0.188 0.290
SigLIP2 [54] ViT-SO400M-16-384 0.176 0.272
DFN [15] ViT-H-14-378 0.220 0.347
PE-core [6] BigG-14-448 0.207 0.312
Analysis using null baseline for steering scores. In the previous analysis of interpretability vs steerability in SAEs, we used the average (or mean) score as the threshold for both metrics. Another choice could be a null baseline for steerability, which would be the average steerability score of the original model’s neurons without an SAE. We found this steerability score of CLIP layer 22 neurons (which are the input to LLaVA SAE/CB-SAE) to be 0.173, while the average steerability score in Fig. 3 (main paper) was 0.232. Using 0.173 as the steerability threshold, the proportions of neurons change as follows:
  • LLaVA (CLIP-ViT-L + Vicuna-7B, Fig. 3, main paper):
    • $\circ$
      Low interp., low steer.: 36.26% $\to$ 17.86%
    • $\circ$
      High interp., low steer.: 19.87% $\to$ 9.58%
    • $\circ$
      Low interp., high steer.: 25.03% $\to$ 43.42%
    • $\circ$
      High interp., high steer.: 18.84% $\to$ 29.14%
However, our original claim of SAEs having low proportion (29%) of high utility neurons is still valid.
Traditional SAE metrics for 4 neuron classes. Reconstruction loss $v-\hat{v}||_{2}^{2}$ (Sec. 3, main paper) does not directly involve SAE latent neurons. To isolate and quantify their impact, we individually zero each SAE neuron and compute the reconstruction loss change w.r.t. the original SAE. Then, the average loss changes over Interpretability/Steerability neuron classes (from Fig. 3) are: low/low (1.6e-5), high/low (1.4e-6), low/high (1.1e-5), high/high (3.9e-6). We observe low interpretability neurons contribute more to reconstruction, while steerability is relatively less influential.
C.3 Extended Analysis of our CB-SAE
Sensitivity to type of SAE. In Table 3, we evaluate the sensitivity of our CB-SAE to the type of pretrained SAE used. We consider Top-$k$ and Batch Top-$k$ SAEs in addition to the Matryoshka SAEs already compared in the main paper. We find that our CB-SAE can provide consistent improvements regardless of the type of pretrained SAE used. Interestingly, Top-$k$ and Batch Top-$k$ SAEs have better interpretability and steerability than Matryoshka SAEs, but also feature a very high number of dead neurons, i.e. SAE neurons which do not activate for any inputs. This makes sense since Matryoshka SAEs were proposed to overcome the dead neurons limitation. Further, our CB-SAE can also resolve the dead neurons problem by eliminating frequently activating SAE neurons.
Ablation of SAE from CB-SAE. In Table 4, we evaluate a CB-SAE model without using any SAE, i.e. a CB-AE [25] where all the user-defined concepts are directly used in the concept bottleneck. We find that CB-AE has higher interpretability score than CB-SAE since all concepts are now explicitly optimized for it. On the other hand, CB-SAE achieves better steerability since the retained SAE neurons have high steerability based on our analysis.
@@ C.3 Extended Analysis of our CB-SAE
Sensitivity to scores used for SAE pruning. We extend our sensitivity analysis from Fig. 5A (main paper) in Table 5 to additionally include monosemanticity score [37] (interpretability evaluation) and zero-shot ImageNet-1k accuracy (reconstruction evaluation) when using the SAE/CB-SAE reconstructed latents. We observe that using either the interpretability score or both scores yields similar reconstruction as the baseline SAE, while steerability-based pruning leads to significantly worse reconstruction. Similarly, using either the interpretability score or both scores improves the monosemanticity significantly w.r.t. the baseline, while steerability-based pruning provides only a marginal gain over the baseline.
@@
Sensitivity to CLIP model in interpretability evaluation. We evaluate the sensitivity of our interpretability evaluation with CLIP-Dissect by varying the CLIP-like model used, in Table 6. While our evaluation used a stronger CLIP-ViT-L-14-336 [42] model w.r.t. the smaller CLIP-ViT-B-16 used for training the CB-SAE, we now evaluate with even stronger models including SigLIP [64], SigLIP2 [54], Data Filtering Networks (DFN) [15] and Perception Encoder (PE) [6]. Across all CLIP-like models, our CB-SAE achieves consistent gains over the baseline SAE for LLaVA with CLIP-ViT-L encoder, validating that our choice of CLIP-like model for interpretability score does not affect our evaluation.
@@
Correlation between LLaVA and UnCLIP steerability. We used the same training method for CB-SAEs in both LLaVA and UnCLIP, but existing pretrained LLaVA and UnCLIP models do not use the same CLIP model (and even use different layers of CLIP and different number of tokens). However, we computed the Pearson correlation between the steerability scores of 1329 common CB concepts from LLaVA and UnCLIP CB-SAEs, which is 0.06 with a p-value 0.035. Hence, steerability seems to be largely downstream task/model-dependent.
@@
C.4 Extended Qualitative Results